INAF queue observing, 2018 Jan 26-27

We (the observers) take this space to bring attention of the world to the the awkward situation of large part of research personnel in INAF. We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world. But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.

However, we believe that like for many other INAF post-docs, our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing. Like many other post-docs, we are paid on contracts renewed every year pending the availability of external funds.

Moreover, we are not entitled to officially say a word on institutional choice and in many aspects we are not considered as INAF employees, but like external collaborators. Indeed, we cannot even lead INAF projects and build a independent research.

Now, we believe INAF has given the power and instruments to start changing this situation.

Observer: R. Carini, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCI2

Summary

Night shared with LINC team.

At the midnight sky is mostly clear, seeing variable
instruments: LBC works, LUCI1 only imaging, LUCI2 imaging and LS, AO works just for LUCI1, MODS1 works, MODS2 only blue channel.

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Night shared with LINC team. INAF time 8:30pm- 13:15pm UT = 4 h 45m

Observing time: 4 h 45m (100%)

Weather time loss: 0 h (0%)

technical time loss 0 h (0%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Piconcelli A2 J0818+0958 LUCI2 HKspec 0.8 h completed
telluric A2 HIP43018 HKspec    
Piconcelli A7 J1210+1741 LUCI2 HKspec 0.3 h completed
telluric A3 HIP54608 LUCI2 HKspec    
Piconcelli A3 J1025+2454 LUCI2 HKspec 1.3 h+ 0.5 completed but to be checked due bad seeing

Details

08:30 we are ready, start with Piconcelli program, the seeing is variable between 1,0" and 1.5"
  • 08:40 send the preset for Piconcelli_A2
  • acq luci2.20180127.0001-0004
  • sci luci2.20180127.0005-0016
  • 08:55 seeing 0.9" but variable
  • 09:14 seeing variable, between 0.8'' and 1.6'' (is due to the passage of warm air)

09:50 we slew to telluric for Piconcelli HIP 43018
  • acq luci2.20180127.0017-0019
  • sci luci2.20180127.0020-0021

10:00 we slew to Piconcelli A7
  • seeing variable, between 1'' and 1.5''
  • acq luci2.20180127.0022-0024
  • sci luci2.20180127.0025-0029
  • seeing 0.8''0.9'''

10:41 we slew to Gilli program, target J
  • acq luci2.0030-0031
  • 10:50 the acquisition is not possible, there is a saturated star near the reference one.

10:51 since the seeing is again variable, up to 1.5'', we slew to Piconcelli program target A3, we begin with the telluric HIP54608
  • acq ==luci2.0032-0036
  • sci luci2.0037-0038

11:13 load A3 scientific target J1025+2454
  • acq luci2.0039-0042
  • sci: luci2.0043-0060
  • seeing 1.4''
  • 11:36 seeing 2.5''
  • 11:41 seeing 1.3''-1.4''
  • 11:56 seeing 2.2''
  • 12:31 seeing 2.0''
  • 12:29 seeing 1.8''
  • 12:35 seeing 1.5''
  • 12:40 seeing 1.3'' variable

12:50 since we have taken some spectra with seeing out of spec, we load again the script to take other spectra
  • seeing 1.3''
  • luci2.0061-0067
  • seeing 1.9''
  • 13:13 seeing 1.5''-1.6'' variable

-- %USERSIG{AndreaRossi - 2018-01-27}%

Comments

 
Topic revision: r12 - 01 Feb 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback